OpenStudy (anonymous):

the electric force between two charged objects is 0.24 newtons. If the distance between them is increased by a factor of 4, what will be the new force? Select one of the options below as your answer: A. 0.03 newtons B. 0.06 newtons C. 0.12 newtons D. 0.36 newtons E. 0.96 newtons

OpenStudy (souvik):

the force is inversely proportional to distance squared

OpenStudy (souvik):

okay... \[F=k/r ^{2}\] here F=force ,k= some constant and r=distance

OpenStudy (anonymous):

0.24=k/4^2

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

nvm cant be

OpenStudy (anonymous):

?

OpenStudy (souvik):

no...you have two situation... \[F=k/r^2\] here F=0.24 N so k/r^2=0.24 now F'=k/(4r)^2 or F'=k/r^2*1/16 or F'=0.24/16

OpenStudy (anonymous):

o.o15?

OpenStudy (anonymous):

can u tell me the answer im seriously feeling super dumb right now

OpenStudy (anonymous):

thanks for the help